A sample of C5+ ions is prepared with the electron in the n = 7 state. The system emits a photon with a wavelength of 343.7 nm. Which shell (value of n) does the electron end up in?

Please enter your answer as a numeric integer. For example, if your answer is one, enter 1
Answer #1

wavelength = 343.7 nm

= 3.437*10^-7 m

Here photon will be emitted

1/lambda = R*Z^2* (1/nf^2 - 1/ni^2)

R is Rydberg constant. R = 1.097*10^7

Here Z = 6 for carbon

1/lambda = R*Z^2* (1/nf^2 - 1/ni^2)

1/3.437*10^-7 = - 1.097*10^7* 6^2*(1/nf^2 - 1/7^2)

(1/nf^2 - 1/7^2) = 7.367*10^-3

1/nf^2 = 2.778*10^-2

nf^2 = 36

nf = 6

Answer: 6